show that the points (-2, 5), (-2, -1) and (4, -1) all lie on a circle whose is at (1, 2). find the length of the radius.
Answer by richwmiller(17219)

(x-4)^2+(y+1)^2=z^2,
(x+2)^2+(y-5)^2=z^2,
(x+2)^2+(y+1)^2=z^2
(x-1)^2+(y-2)^2 = 18
radius = 3 sqrt(2)=4.24264